An object starts from rest at time t = 0.00 s and moves in the +x direction with constant acceleration. The object travels 12.0 m from time t = 1.00 s to time t = 2.00 s. Calculate the acceleration of the object.


1
Expert's answer
2021-06-18T11:21:49-0400

"x(t) = v_0t +\frac{at^2}{2}"

"v_0 =0"

"x(t) = \frac{at^2}{2}"

"x(1) = \frac{a}{2}"

"x(2) = 2a"

"x(2)-x(1)=12"

"x(2)-x(1)=2a-\frac{a}{2}=\frac{3a}{2}"

"\frac{3a}{2}=12"

"a=8"

"\text{Answer: }8 \ \frac{m}{s^2}"